Exercise 11-06 Your answer is partially correct. Try again. Metlock Company purchased equipment for $285,600 on October 1, 20Just need help with d) and e). The answer for d) is NOT 45600 or 11400. The answer for e) is NOT 71400 or 53550.
Answer #1

Answers

  • Requirement a,b and c are ALREADY CORRECR
  • Requirement [d] = $ 51,300

Working

A

Cost

$          285,600.00

B

Residual Value

$            12,000.00

C=A - B

Depreciable base

$          273,600.00

Year

Depreciable base

Formula

Depreciation expense

2020

$           273,600.00

[273600 x 8/36 x 3months/12months]

$         15,200.00

2021

$           273,600.00

[273600 x 8/36 x 9months/12months] + [273600 x 7/36 x 3/12]

$         58,900.00

2022

$           273,600.00

[273600 x 7/36 x 9months/12months] + [273600 x 6/36 x 3/12]

$         51,300.00

  • Requirement [e] = $ 66,938

Working

A

Cost

$          285,600.00

B

Residual Value

$            12,000.00

C=A - B

Depreciable base

$          273,600.00

D

Life [in years]

8

E=C/D

Annual SLM depreciation

$            34,200.00

F=E/C

SLM Rate

12.50%

G=F x 2

DDB Rate

25.00%

Year

Beginning Book Value

Depreciation rate

Depreciation expense

Ending Book Value

2020

$           285,600.00

25.00%

$         17,850.00 [$ 285600 x 25% x 3/12]

$           267,750.00

2021

$           267,750.00

25.00%

$         66,937.50

$           200,812.50